# Neural Isometries: Taming Transformations for Equivariant ML

### v1

@@ B.3 Homography-Perturbed MNIST
The autoencoders are trained for $50,000$ steps with a batch size of $16$ using the AdamW optimizer with a weight decay of $10^{-4}$. The learning rate follows a schedule consisting of a 2,000 step warm up from $0.0$ to $5\times 10^{-4}$ and afterwards decays to $5\times 10^{-5}$ via cosine annealing. Training takes approximately 30 minutes.
@@
As no code release exists for the NFT, we implement it following the procedure described in the paper [35]. While a ViT-based encoder is proposed in the original paper, we find that it produces worse evaluation results than the 2-layer ConvNet encoder described above and thus use the latter in our evaluations to provide a fair comparison. We follow the authors’ proposed implementation and consider only a reconstruction loss and learn a diagonalization transform in a post-processing step. Otherwise, the NFT is pre-trained, fine-tuned, and evaluated identically to our own approach, and uses the same equivariant classification head. We note that the dimension of the latent map constructed by the NFT is the same as $\tau{}_{\Omega{}}{}$ $(32\times 32)$ and that the tensor passed to the classification head is the same size for both our method and the NFT.
The autoencoder baseline consists of the same encoder and decoder, pre-trained identically but with respect to the standard reconstruction loss $\lVert\mathcal{D}{}(\mathcal{E}{}(\psi))-\psi\rVert$ without considering $T$-related pairs. During the fine tuning phase, latents are passed directly to a standard non-equivariant 2-layer MLP with 128 channels to form class predictions. Otherwise, fine-tuning and evaluation are performed in the same manner.
@@ B.4 Conformal Shape Classification
Training is performed with respect to the composite loss with $\alpha=0.5$ and $\beta=0.01$. The model is trained for 50,000 iterations with a batch size of 16 using the AdamW optimizer with a weight decay of $10^{-4}$. The learning rate follows a schedule consisting of a 2,000 step warm up from $0.0$ to $5\times 10^{-4}$ and afterwards decays to $5\times 10^{-5}$ via cosine annealing. Training takes approximately four hours.
@@
We pre-train, fine-tune, and evaluate the NFT and autoencoder baseline using the same autoencoder architecture, classification head, and training regimes as our own method in the same manner as sec. (B.3). Again, the both the size of the latent map and size of the tensor passed to the classification are the same size for both our method and the NFT. Here, the weights of the encoder are left unfrozen, and are jointly optimized with the classification head during the fine-tuning phase.
@@ B.5 Camera Pose Estimation from Real-World Video
The prediction head is trained for 20,000 iterations with a batch size of 16 (consisting of 4 sub-batches of pairs of frames from 4 sequences) using the AdamW optimizer with the same weight decay and learning rate schedule as used previously. Training takes approximately one 80 minutes.
@@
The transformer baseline is based on the CroCo/DUSt3R[46, 47] architectures. As with our approach, input frames are passed to the same decoder. Here the encoder consists of 12-layer VITs with a patch size of 16, with 16 heads and 512 channels per layer. Adjacent frames are independently encoded, and then concatenated along the spatial dimension in addition to a single learned token. The resulting tensor is passed to a 12-layer transformer decoder which mirrors the encoder with 16 heads and 512 channels per layer. Subsequently, the learned token is extracted from the output and passed through a linear layer to recover the predicted pose parameters. The model is trained from scratch in the fine-tuning phase, using the same loss and optimization procedure as described above.
@@ Appendix C Additional Results
Below we provide additional qualitative results including visualizations of the mass matrices $\mathbb{M}{}$ learned in each experiment and additional comparisons of estimated camera trajectories recovered by each method in the pose prediction experiments. For the latter, examples of failure cases are also included.
Figure 4: Visualizing the Mass Matrix $\mathbb{M}$ Elements at Grid Indices. Visualizations of the elements of the diagonal mass matrices $\mathbb{M}$ learned in each experiment, in terms of the magnitude of the deviation from the mean value at each grid index in the latent space. White indicates little deviation from the mean, with green-blue indicating mass values above the mean and orange-red indicating mass values below. When approximating the toric Laplacian (sec. 4.3), the learned mass matrix is approximately constant, mirroring the standard geometry of the toric grid, which assumes unit mass at each index. In the MNIST experiments (sec. 5.1), the distribution of mass appears to segment null space from the central region most often occupied by the digits. In the conformal shape classification experiments (sec. 5.2), the larger deviations from the mean values appear closer to the poles (the top-most and bottom-most rows of the spherical grid). For the post estimation experiments (sec. 5.3), larger deviations appear at the boundaries, with the lower half of the grid having slightly higher values.

### v2

@@ B.3 Homography-Perturbed MNIST
The autoencoders are trained without spectral dropout for $50,000$ steps with a batch size of $16$ using the AdamW optimizer with a weight decay of $10^{-4}$. The learning rate follows a schedule consisting of a 2,000 step warm up from $0.0$ to $5\times 10^{-4}$ and afterwards decays to $5\times 10^{-5}$ via cosine annealing. Training takes approximately 30 minutes.
@@
We implement the NFT following the procedure described in the paper [37]. While a ViT-based encoder is proposed in the original paper, we find that it produces worse evaluation results than the 2-layer ConvNet encoder described above and thus use the latter in our evaluations to provide a fair comparison. We follow the authors’ proposed implementation and consider only a reconstruction loss and learn a diagonalization transform in a post-processing step. Otherwise, the NFT is pre-trained, fine-tuned, and evaluated identically to our own approach, and uses the same equivariant classification head. We note that the dimension of the latent map constructed by the NFT is the same as $\tau{}_{\Omega{}}{}$ $(32\times 32)$ and that the tensor passed to the classification head is the same size for both our method and the NFT.
The autoencoder baseline consists of the same encoder and decoder, pre-trained identically but with respect to the standard reconstruction loss $\lVert\mathcal{D}{}(\mathcal{E}{}(\psi))-\psi\rVert$ without considering $T$-related pairs. During the fine-tuning phase, we apply data augmentation to the input images by transforming them by homographies sampled from the distribution proposed in [6] before encoding. Latents are passed directly to a standard non-equivariant 2-layer MLP with 128 channels to form class predictions. Otherwise, fine-tuning and evaluation are performed in the same manner.
@@ B.4 Conformal Shape Classification
Training is performed with respect to the composite loss with $\alpha=0.5$ and $\beta=0.01$. The model is trained without spectral dropout for 50,000 iterations with a batch size of 16 using the AdamW optimizer with a weight decay of $10^{-4}$. The learning rate follows a schedule consisting of a 2,000 step warm up from $0.0$ to $5\times 10^{-4}$ and afterwards decays to $5\times 10^{-5}$ via cosine annealing. Training takes approximately four hours.
@@
We pre-train, fine-tune, and evaluate the NFT and autoencoder baseline using the same autoencoder architecture, classification head, and training regimes in the same manner as sec. (B.3). Here data augmentation is applied during the fine-tuning phase by randomly sampling and applying Möbius transformations with scale factors proportional to those estimated between the spherical parameterizations of the shapes. Again, the both the size of the latent map and size of the tensor passed to the classification are the same size for both our method and the NFT. Here, the weights of the encoder are left unfrozen, and are jointly optimized with the classification head during the fine-tuning phase.
@@ B.5 Camera Pose Estimation from Real-World Video
The prediction head is trained without spectral dropout for 20,000 iterations with a batch size of 16 (consisting of 4 sub-batches of pairs of frames from 4 sequences) using the AdamW optimizer with the same weight decay and learning rate schedule as used previously. Training takes approximately one 80 minutes.
@@
The transformer baseline is based on the CroCo/DUSt3R[50, 51] architectures. As with our approach, input frames are passed to the same decoder. Here the encoder consists of 12-layer VITs with a patch size of 16, with 16 heads and 512 channels per layer. Adjacent frames are independently encoded, and then concatenated along the spatial dimension in addition to a single learned token. The resulting tensor is passed to a 12-layer transformer decoder which mirrors the encoder with 16 heads and 512 channels per layer. Subsequently, the learned token is extracted from the output and passed through a linear layer to recover the predicted pose parameters.
Additionally, we also evaluate the efficacy of DINOv2[47] and BeIT[48] as representation learners for the pose estimation task. Pairs of adjacent frames are passed through the pre-trained models and the feature tokens for each image from the final layers are extracted. These are used as input to the same DUSt3R[47]-style decoder described above which is trained to predict the parameters of the relative camera pose.
The VIT model and DINOv2/BeIT prediction heads are trained from scratch in the fine-tuning phase, using the same loss and optimization procedure as described above.
@@ Appendix C Additional Results
Below we provide additional results including a quantiative evaluation of learned equivariance, visualizations of the eigenfunctions $\Phi$ and mass matrices $\mathbb{M}{}$ learned in each experiment, and comparisons of estimated camera trajectories recovered by each method in the pose prediction experiments. For the latter, examples of failure cases are also included.
CO3D Equivariance Error
Frame Skip 0 1 3 5 7 9
NIso 5.25% ($\pm$ 1.61) 5.84% ($\pm$ 1.41) 7.44% ($\pm$ 1.43) 8.23% ($\pm$ 1.48) 8.79 % ($\pm$ 1.56) 9.24% ($\pm$ 1.67)
Equivariance Error
Hom. MNIST Conf. SHREC ‘11
NIso 4.06% ($\pm$ 3.79) 5.91% ($\pm$ 3.49)
homConv [6] 6.86%
MobiusConv [7] 2.64%
Table 4: Quantitative Evaluation of Learned Equivariance. Mean NIso equivariance error in the latent space across all experiments. Following [6, 7, 45], we measure the learned equivariance of the NIso latent space in the standard way via the average of $\lVert\tau{}_{\Omega{}}{}\,\mathcal{E}{}_{\Omega{}}(\psi)-\mathcal{E}{}_{\Omega{}}(T\psi)\rVert^{2}/\lVert\mathcal{E}{}_{\Omega{}}(T\psi)\rVert^{2}$ over all pairs in the test set, for five randomly initialized pre-training runs. For Hom. MNIST, we randomly sample homographies from the distribution proposed in [6] which are applied to the elements of the standard MNIST test set; for Conf. SHREC ‘11, pairs are formed from the sets of conformally- augmented meshes derived from the same base shape in the test split. With CO3D we measure the error between encodings of adjacent frames in test set, with increasing frame skip. We also list the errors reported by competing hand-crafted methods.
Figure 4: Visualizing the Learned Eigenfunctions $\Phi$ and Mass Matrices $\mathbb{M}{}$. Visualizations of the eigenfunctions $\Phi$ learned in each experiment are shown on the top row. Eigenfunctions are sorted by eigenvalue in ascending order along rows in C-style indexing. Here, experiments were performed without spectral dropout so the ordering is random. The elements of the learned diagonal mass matrices $\mathbb{M}$ are shown on the bottom row, in terms of the magnitude of the deviation from the mean value at each grid index in the latent space. White indicates little deviation from the mean, with green-blue indicating mass values above the mean and orange-red indicating mass values below. In the MNIST experiments (sec. 5.1), the distribution of mass appears to segment null space from the central region most often occupied by the digits. In the conformal shape classification experiments (sec. 5.2), the larger deviations from the mean values appear closer to the poles (the top-most and bottom-most rows of the spherical grid). For the pose estimation experiments (sec. 5.3), larger deviations appear at the boundaries, with the lower half of the grid having slightly higher values.